OpenStudy (anonymous):

easy questions i think If I know the lengths of two sides of a right triangle, how do I find the third?

OpenStudy (anonymous):

if it is a right triangle you use the pythagorean... so a^2 + b^2 = c^2. c would be the hypotenuse while a and b are the legs of the triangle.

OpenStudy (anonymous):

Use the formula a^2 + b^2 = c^2 c being the hypontenus and a and b being the other sides.